OpenStudy (anonymous):

A company distributes free erasers to all the students of x schools. Each school has (x + 2) classes. The number of students in each class is 1 less than the number of classes in each school. Each student is given (x - 2) erasers. Part A: Write an expression to show the total number of erasers distributed by the company in x schools. (4 points) Part B: What would x(x + 2) represent? When simplified, what would be the DEGREE and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points) Can someone help me?

OpenStudy (anonymous):

I think that part A is: (x+2)-1(x-2) B: x^2 +2x and degree of 3 and it is a binomial I am not sure if I am right

OpenStudy (anonymous):

Each school has (x + 2) classes

OpenStudy (anonymous):

The number of students in each class is 1 less than the number of classes in each school that makes the number of students in each class \(x+2-1=x+1\)

OpenStudy (anonymous):

there are \(x+1\) students each given \(x-2\) erasers

OpenStudy (anonymous):

that makes the total \((x+1)(x-2)\)

OpenStudy (anonymous):

B is right

OpenStudy (anonymous):

So to find the total number of students I just combine?

OpenStudy (anonymous):

x^2 -2x +1x -2

OpenStudy (anonymous):

x^2 - x - 2

OpenStudy (anonymous):

actually i have no idea how to calculate the number of students

OpenStudy (anonymous):

ok. thank you for your help

OpenStudy (anonymous):

oh yes i do!!

OpenStudy (anonymous):

(x + 2) classes. The number of students in each class is 1 less than the number of classes in each school.

OpenStudy (anonymous):

that makes the total number of students \((x+2)(x+1)=x^2+3x+2\)